A barge traveling down a river covers a distance of 36 miles in 4 hours. It would have traveled the same distance in 6 hours if it had been traveling upstream against the current. What is the speed of the river current?
Answer by lwsshak3(11628)

A barge traveling down a river covers a distance of 36 miles in 4 hours. It would have traveled the same distance in 6 hours if it had been traveling upstream against the current. What is the speed of the river current?
***
let x=speed of barge in still water
x+c=speed of boat downstream
x-c=speed of boat upstream
travel time*speed=distance
..
4(x+c)=36
6(x-c)=36
..
4x+4c=36
6x-6c=36
..
12x+12c=108 (mult. by 3)
12x-12c=72 (mult. by 2)
subtract:
24c=36
c=36/24=3/2=1.5
speed of the river current=1.5 mph